An electric field of magnitude 11000 N/C

and directed upward (perpendicular to the

Earth
Answer #1

Electric flux = E.A = E * A * cos 86 = 11000 * 8.8 * 3.5 cos 86 = 23633.50 Nm^2/C
Answer #2

Electric flux

= E.A

= E * A * cos 86

= 11000 * 8.8 * 3.5 cos 86

= 23633.50 Nm^2/C